### v1

# Characterizing Kadison–Schwarz maps on $M_{3}$

Abstract
Kadison–Schwarz (KS) maps form a natural class of positive linear maps lying between positivity and complete positivity. Despite their relevance in quantum dynamics and operator algebras, a detailed characterization of KS maps beyond low dimensions remains largely open. In this work we analyze unital linear maps on $M_{3}$ using the Bloch–Gell–Mann representation. Exploiting unitary equivalence and structural properties of the $\mathfrak{su}(3)$ algebra, we derive analytic conditions ensuring the Kadison–Schwarz property. Our approach clarifies the relation between KS maps and completely positive maps on $M_{3}$.
I Introduction
Positive linear maps between matrix algebras play a fundamental role in quantum information theory, operator algebras, and the theory of open quantum systems. Among various classes of positive maps, completely positive (CP) maps occupy a distinguished position due to their physical interpretation as quantum channels [Choi1975]. However, in several contexts—notably in the analysis of quantum dynamical maps, divisibility, and non-Markovian dynamics—one encounters broader classes of positive maps for which complete positivity is not required, see, e.g., Refs. [Chruscinski2010, Chruscinski2017].
An important example of such a class is formed by Kadison–Schwarz (KS) maps. These maps are characterized by a quadratic operator inequality and occupy an intermediate position between positivity and complete positivity. While every completely positive map satisfies the Kadison–Schwarz property, the converse implication is false. As a result, KS maps constitute a genuinely intermediate class of positive maps with properties that are neither purely algebraic nor fully captured by complete positivity.
Several general features of KS maps are known. In particular, unital positive maps satisfy the Kadison–Schwarz inequality when restricted to normal operators, whereas explicit counterexamples show that unitality and positivity alone do not guarantee the KS property. Beyond such general observations, however, explicit analytic characterizations of KS maps remain scarce and are typically restricted to low-dimensional or highly symmetric settings.
The aim of the present work is to analyze Kadison–Schwarz maps within the Bloch–Gell–Mann framework. This representation provides a natural parametrization of linear maps on $M_{d}$ adapted to the $\mathfrak{su}(d)$ structure and allows one to systematically exploit unitary equivalence and canonical decompositions of the associated Bloch matrix. While the framework itself applies to arbitrary dimension $d$, the derivation of explicit analytic conditions depends sensitively on the underlying algebraic structure.
In this work we derive explicit analytic sufficient conditions for the Kadison–Schwarz property in the case of the matrix algebra $M_{3}$. The special role of $M_{3}$ stems from the tractable structure of $\mathfrak{su}(3)$, which permits nontrivial analytic bounds without resorting to numerical optimization. Our results clarify the relation between Kadison–Schwarz maps and completely positive maps on $M_{3}$ and identify structural features of the Bloch–Gell–Mann representation that are responsible for the KS property. To the best of our knowledge, explicit analytic sufficient conditions for the Kadison–Schwarz property of unital maps on $M_{3}$, expressed directly in terms of Bloch (Gell–Mann) parameters, have not been systematically discussed in the literature beyond the qubit case ($d=2$) or highly symmetric subclasses.
The paper is organized as follows. In Sec. II we recall the definition of Kadison–Schwarz maps and review their basic properties. In Sec. III we introduce the Bloch–Gell–Mann representation of unital maps on $M_{d}$. Our main results, specialized to the case of $M_{3}$, are presented in Sec. IV. Technical derivations are collected in the Appendix.
II Kadison–Schwarz maps
Let $\Phi:M_{d}\to M_{d}$ be a linear map. We say that $\Phi$ is positive if $\Phi(X)\geq 0$ for all $X\geq 0$, and unital if $\Phi(\mathbb{I})=\mathbb{I}$. The Kadison–Schwarz (KS) property is defined by the operator inequality
$\Phi(X^{\dagger}X)\;\geq\;\Phi(X^{\dagger})\,\Phi(X),$ (1)
which is required to hold for all $X\in M_{d}$. The inequality (1) was originally introduced by Kadison in the context of operator algebras [Kadison1952]. We emphasize that the KS property does not, in general, imply unitality, although the two notions are often considered together.
Every completely positive (CP) map satisfies the Kadison–Schwarz inequality, but the converse implication does not hold. Hence KS maps form a strictly larger class than CP maps and interpolate between complete positivity and positivity,
$\mathrm{CP}\;\subsetneq\;\mathrm{KS}\;\subsetneq\;\mathrm{Pos},$ (2)
see, e.g., Refs. [Chruscinski2010, Mukhamedov2011]. Despite this simple ordering, the structure of KS maps is considerably less understood than that of completely positive maps.
We next recall several basic properties of unital positive maps that will be useful in the analysis of the Kadison–Schwarz condition, starting with a classical result due to Kadison.
Lemma 1 (Kadison inequality).
Let $\Phi:M_{d}\to M_{d}$ be a unital positive map. Then, for every Hermitian operator $A\in M_{d}$,
$\|\Phi(A)\|\;\leq\;\|A\|,$ (3)
where $\|\cdot\|$ denotes the operator norm.
Lemma 1 implies that unital positive maps are contractions on the space of Hermitian operators. Importantly, this property does not rely on complete positivity and plays a key role in the analysis of KS maps [Kadison1952, Chruscinski2010].
The following lemma clarifies to what extent the KS inequality is automatically satisfied by unital positive maps.
Lemma 2.
Let $\Phi:M_{d}\to M_{d}$ be a unital positive map. Then the Kadison–Schwarz inequality (1) holds for all normal operators $X\in M_{d}$, i.e., for operators satisfying $X^{\dagger}X=XX^{\dagger}$.
Proof.
Let $X$ be a normal operator. Then $X^{\dagger}X=XX^{\dagger}\geq 0$ and $\|X\|^{2}=\|X^{\dagger}X\|$. By positivity, $\Phi(X^{\dagger}X)\geq 0$. Using Lemma 1, we obtain
$\|\Phi(X)\|^{2}\;\leq\;\|X\|^{2}\;=\;\|X^{\dagger}X\|.$
For any operator $B$, one has $B^{\dagger}B\leq\|B\|^{2}\mathbb{I}$, and hence
$\Phi(X)^{\dagger}\Phi(X)\;\leq\;\|\Phi(X)\|^{2}\mathbb{I}\;\leq\;\|X^{\dagger}X\|\,\mathbb{I}.$
Since both $\Phi(X^{\dagger}X)$ and $\Phi(X^{\dagger})\Phi(X)$ are positive operators bounded above by the same scalar multiple of the identity, the inequality (1) follows. ∎
Lemma 2 shows that the KS property does not impose additional constraints on unital positive maps when restricted to normal operators. Consequently, possible violations of the KS inequality must originate from non-normal operators, cf. [Chruscinski2010, Mukhamedov2011].
We now recall a standard example illustrating that unital positive maps need not satisfy the Kadison–Schwarz inequality.
The transposition map $T:M_{d}\to M_{d}$, defined by $T(X)=X^{T}$, is positive and unital, but it does not satisfy the Kadison–Schwarz property. Indeed, for a rank-one operator $X=|1\rangle\langle 2|$ one has
$X^{\dagger}X=|2\rangle\langle 2|,\qquad XX^{\dagger}=|1\rangle\langle 1|,$
which leads to
$T(X^{\dagger}X)=|2\rangle\langle 2|,\qquad T(X^{\dagger})\,T(X)=|1\rangle\langle 1|.$
Hence the inequality $T(X^{\dagger}X)\geq T(X^{\dagger})T(X)\;$ is violated. This example is well known and discussed in detail in the literature, see, e.g., Refs. [Chruscinski2010, Mukhamedov2011].
The transposition map therefore provides a simple illustration that the Kadison–Schwarz condition is genuinely stronger than positivity, even in the unital case, and motivates the search for explicit criteria distinguishing KS maps from merely positive ones.
III Bloch–Gell–Mann representation of unital maps on $M_{d}$
We emphasize that the Bloch representation of linear maps does not depend on the particular choice of an orthonormal basis of the Lie algebra $\mathfrak{su}(d)$. Different choices, such as the Weyl or generalized Pauli bases, are related by orthogonal transformations in the space of Bloch vectors and therefore lead to equivalent descriptions of unital maps [Kimura2003, BengtssonZyczkowski]. In particular, the Kadison–Schwarz property is invariant under such basis changes. The choice of the Gell–Mann basis is motivated by technical convenience: its generators are Hermitian and possess well-known symmetric and antisymmetric structure constants, which allow for a transparent algebraic treatment of operator products appearing in the Kadison–Schwarz inequality [Georgi1999]. This makes the Gell–Mann representation particularly suitable for deriving analytic bounds.
III.1 Bloch representation on $M_{d}$
Let $\{\lambda_{k}\}_{k=1}^{d^{2}-1}$ be a fixed orthonormal basis of $\mathfrak{su}(d)$ consisting of Hermitian, traceless matrices, normalized as
$\operatorname{Tr}(\lambda_{i}\lambda_{j})=2\,\delta_{ij}.$ (4)
The symmetric and antisymmetric structure constants $d_{ijk}$ and $f_{ijk}$ are defined in the standard way via the anticommutation and commutation relations of the generators, respectively. Together with $\lambda_{0}:=\sqrt{\tfrac{2}{d}}\,\mathbb{I}$, this basis provides an orthonormal basis of $M_{d}$ with respect to the Hilbert–Schmidt inner product.
Any operator $X\in M_{d}$ can be written in the Bloch form
$X=w_{0}\lambda_{0}+\sum_{k=1}^{d^{2}-1}w_{k}\lambda_{k},$ (5)
where $w_{0}\in\mathbb{R}$ and $w_{k}\in\mathbb{C}$. For Hermitian operators $X$, all coefficients $w_{k}$ are real [Kimura2003, BengtssonZyczkowski].
The product of two basis elements can be expressed as
$\lambda_{i}\lambda_{j}=\frac{2}{d}\,\delta_{ij}\,\mathbb{I}+\sum_{k=1}^{d^{2}-1}\left(d_{ijk}+if_{ijk}\right)\lambda_{k},$ (6)
where $f_{ijk}$ and $d_{ijk}$ are the completely antisymmetric and symmetric structure constants of $\mathfrak{su}(d)$, respectively [Georgi1999, Slansky].
III.2 Unital maps in the Bloch representation
Let $\Phi:M_{d}\to M_{d}$ be a unital linear map. In the Bloch representation (5), such a map acts affinely as
$\Phi(X)=w_{0}\lambda_{0}+\sum_{k=1}^{d^{2}-1}(Tw)_{k}\lambda_{k},$ (7)
where $w=(w_{1},\dots,w_{d^{2}-1})^{T}$ and $T$ is a real $(d^{2}-1)\times(d^{2}-1)$ matrix, referred to as the Bloch matrix of $\Phi$ [Kimura2003, Ruskai2002]. Unitality of $\Phi$ implies that the identity component $w_{0}$ is preserved, while positivity imposes nontrivial constraints on $T$.
The Bloch matrix representation is not unique. If $\Phi$ is conjugated by unitary operators,
$\Phi\;\mapsto\;\mathcal{U}_{1}\circ\Phi\circ\mathcal{U}_{2},\qquad\mathcal{U}_{i}(X)=U_{i}XU_{i}^{\dagger},$ (8)
then the corresponding Bloch matrix transforms as
$T\;\mapsto\;O_{1}\,T\,O_{2},$ (9)
where $O_{1},O_{2}\in SO(d^{2}-1)$ are orthogonal matrices induced by the adjoint action of $SU(d)$ on $\mathfrak{su}(d)$ [BengtssonZyczkowski].
III.3 Canonical form
Using the polar decomposition, any Bloch matrix $T$ can be written as
$T=RS,$ (10)
where $R\in SO(d^{2}-1)$ is orthogonal and $S$ is symmetric and positive semidefinite. Since orthogonal transformations correspond to unitary conjugations, which leave the Kadison–Schwarz property invariant, one may restrict attention to Bloch matrices of the form $T=S$ [Ruskai2002].
Consequently, without loss of generality, unital maps can be represented by symmetric Bloch matrices. In particular, if $S$ is diagonalizable, the map $\Phi$ is unitarily equivalent to a map with diagonal Bloch matrix
$T=\operatorname{diag}(\mu_{1},\mu_{2},\dots,\mu_{d^{2}-1}),$ (11)
where the real parameters $\mu_{k}$ characterize the action of $\Phi$ on the traceless subspace.
This canonical form provides a convenient starting point for analyzing positivity and Kadison–Schwarz conditions. In the following section we specialize this framework to the case of $M_{3}$, where the explicit structure of $\mathfrak{su}(3)$ allows for analytic estimates.
IV Main result: Kadison–Schwarz condition on $M_{3}$
In this section we specialize the general Bloch–Gell–Mann framework developed above to the case of the matrix algebra $M_{3}$. The special role of $M_{3}$ stems from the explicit structure of the Lie algebra $\mathfrak{su}(3)$, which allows one to control the structure constants appearing in operator products and to derive analytic estimates for the Kadison–Schwarz inequality.
Theorem 1.
Let $\Phi:M_{3}\to M_{3}$ be a unital linear map with diagonal Bloch matrix
$T=\operatorname{diag}(\mu_{1},\mu_{2},\dots,\mu_{8})$
in the Gell–Mann representation. Assume that
$\max_{1\leq k\leq 8}|\mu_{k}|\leq 1$ (12)
and that the parameters $\{\mu_{k}\}$ satisfy
$\max_{i,j}|\mu_{i}-\mu_{j}|\;\leq\;\frac{2\sqrt{2}}{3\,C_{3}},$ (13)
where $C_{3}$ is a positive constant depending only on the symmetric structure constants $d_{ijk}$ of $\mathfrak{su}(3)$. Then the map $\Phi$ satisfies the Kadison–Schwarz inequality.
Remark 1.
The constant $C_{3}$ arises from norm estimates of expressions involving the symmetric structure constants $d_{ijk}$ of $\mathfrak{su}(3)$. Its explicit value is not essential for the formulation of the result and is discussed in the Appendix. The condition (13) is sufficient but not necessary for the Kadison–Schwarz property.
IV.1 Reduction of the Kadison–Schwarz inequality
Let $X\in M_{3}$ be an arbitrary operator and write it in the Bloch–Gell–Mann form
$X=w_{0}\lambda_{0}+\sum_{k=1}^{8}w_{k}\lambda_{k}.$
Using the product relations for the Gell–Mann generators, the operator $X^{\dagger}X$ can be expressed as a linear combination of the identity and traceless generators. Applying the unital map $\Phi$ with diagonal Bloch matrix $T=\operatorname{diag}(\mu_{1},\dots,\mu_{8})$, the Kadison–Schwarz expression takes the form
$\Phi(X^{\dagger}X)-\Phi(X^{\dagger})\Phi(X)=\alpha(X)\,\mathbb{I}+\sum_{k=1}^{8}\beta_{k}(X)\,\lambda_{k},$ (14)
where the coefficients $\alpha(X)$ and $\beta_{k}(X)$ depend quadratically on the Bloch components $w_{k}$ and on the parameters $\mu_{k}$.
A crucial simplification occurs due to the antisymmetric structure constants $f_{ijk}$ of $\mathfrak{su}(3)$. Since the Bloch matrix $T$ is diagonal and the parameters $\mu_{k}$ are real, all terms proportional to $f_{ijk}$ cancel identically in (14). As a consequence, the nontrivial part of the Kadison–Schwarz inequality is governed solely by contributions involving the symmetric structure constants $d_{ijk}$. This mechanism relies essentially on the presence of nonvanishing symmetric structure constants and therefore does not occur in the qubit case $d=2$, where $d_{ijk}$ vanish identically. It is thus specific to $M_{3}$ and higher-dimensional matrix algebras and plays a key role in enabling analytic estimates.
IV.2 Proof of Theorem 1
By the reduction described above, the Kadison–Schwarz condition is equivalent to the positivity of the operator (14). The identity component $\alpha(X)$ is nonnegative under the contraction assumption (12). The remaining traceless part can be estimated using operator norm bounds for expressions involving the symmetric structure constants $d_{ijk}$.
More precisely, one obtains an inequality of the form
$\biggl\|\sum_{k=1}^{8}\beta_{k}(X)\,\lambda_{k}\biggr\|\;\leq\;C_{3}\,\max_{i,j}|\mu_{i}-\mu_{j}|\,\|X^{\dagger}X\|,$ (15)
where $C_{3}$ depends only on the algebraic structure of $\mathfrak{su}(3)$. The condition (13) guarantees that the traceless contribution in (14) is dominated by the positive identity term. Hence,
$\Phi(X^{\dagger}X)\geq\Phi(X^{\dagger})\Phi(X)$
for all $X\in M_{3}$, which proves the theorem. A detailed derivation of the bound (15) and an explicit expression for the constant $C_{3}$ are provided in the Appendix.
Remark 2.
For completely positive maps the Kadison–Schwarz inequality is automatically satisfied. Theorem 1 shows that, on $M_{3}$, the Kadison–Schwarz property may hold under substantially weaker assumptions, which can be expressed explicitly in terms of the Bloch parameters.
V Conclusions
In this work we analyzed the Kadison–Schwarz (KS) property for unital linear maps acting on matrix algebras, using the Bloch–Gell–Mann representation. After reviewing general properties of KS maps on $M_{d}$, we specialized the analysis to the case of $M_{3}$, where the explicit structure of the Lie algebra $\mathfrak{su}(3)$ allows for analytic control of operator products appearing in the Kadison–Schwarz inequality.
Our main result provides an explicit sufficient condition for the Kadison–Schwarz property of unital maps on $M_{3}$, formulated directly in terms of the parameters of the Bloch matrix. In contrast to approaches based on numerical optimization or on complete positivity criteria, the condition derived here is purely analytic and exploits structural features of the $\mathfrak{su}(3)$ algebra. In particular, the cancellation of all contributions involving the antisymmetric structure constants and the resulting reduction to estimates involving only the symmetric constants play a central role.
The results clarify the position of KS maps between positive and completely positive maps on $M_{3}$. They demonstrate that the Kadison–Schwarz property may hold under assumptions substantially weaker than complete positivity, while still imposing nontrivial constraints beyond mere positivity. This highlights the usefulness of KS maps as an intermediate class, relevant in contexts where complete positivity is not required but additional structure is desirable.
The framework developed in this work naturally extends to arbitrary dimension $d$. The Bloch representation and the reduction via unitary equivalence remain valid on $M_{d}$, and the Kadison–Schwarz property is invariant under changes of the generator basis. However, for $d>3$ the control of expressions involving the symmetric structure constants of $\mathfrak{su}(d)$ becomes increasingly subtle, and the derivation of sharp analytic bounds requires additional tools. A systematic extension of the present analysis to higher dimensions, including dimension-dependent estimates for KS maps on $M_{d}$, is left for future work.
We expect that the methods presented here may also be useful in the study of Kadison–Schwarz properties of dynamical maps and generators appearing in the theory of open quantum systems, as well as in the analysis of intermediate classes of positive maps beyond complete positivity.
Appendix A Estimates involving the symmetric structure constants of $\mathfrak{su}(3)$
In this Appendix we provide the technical estimates underlying Theorem 1. In particular, we derive bounds on expressions involving the symmetric structure constants $d_{ijk}$ of $\mathfrak{su}(3)$ and introduce the constant $C_{3}$ appearing in Eqs. (13) and (15).
A.1 Structure constants of $\mathfrak{su}(3)$
Let $\{\lambda_{k}\}_{k=1}^{8}$ denote the standard Gell–Mann generators, normalized as $\operatorname{Tr}(\lambda_{i}\lambda_{j})=2\delta_{ij}$. Their products satisfy
$\lambda_{i}\lambda_{j}=\frac{2}{3}\,\delta_{ij}\,\mathbb{I}+\sum_{k=1}^{8}\left(d_{ijk}+if_{ijk}\right)\lambda_{k},$ (16)
where $f_{ijk}$ and $d_{ijk}$ are the antisymmetric and symmetric structure constants of $\mathfrak{su}(3)$, respectively. Explicit values of these constants can be found, e.g., in Refs. [Georgi1999, Slansky].
In the analysis of the Kadison–Schwarz inequality only the symmetric constants $d_{ijk}$ contribute. This follows from the cancellation of all terms involving $f_{ijk}$ for diagonal Bloch matrices, as discussed in Sec. IV.
A.2 Norm estimates
Consider expressions of the form
$Y=\sum_{k=1}^{8}y_{k}\lambda_{k},$ (17)
where the coefficients $y_{k}$ arise as quadratic combinations of the Bloch components $w_{i}$ and the parameters $\mu_{j}$. Using the submultiplicativity of the operator norm and the orthonormality of the Gell–Mann basis, one obtains the bound
$\|Y\|\;\leq\;\sum_{k=1}^{8}|y_{k}|\,\|\lambda_{k}\|=\sqrt{2}\,\sum_{k=1}^{8}|y_{k}|.$ (18)
The coefficients $y_{k}$ contain terms of the form
$\sum_{i,j}(\mu_{i}-\mu_{j})\,d_{ijk}\,w_{i}\overline{w_{j}}.$
Applying the Cauchy–Schwarz inequality and using the finiteness of the set of nonvanishing symmetric structure constants, one finds
$\sum_{k=1}^{8}|y_{k}|\;\leq\;C_{3}\,\max_{i,j}|\mu_{i}-\mu_{j}|\,\|X^{\dagger}X\|,$ (19)
where the constant $C_{3}>0$ depends only on the numerical values of the coefficients $d_{ijk}$.
Combining (18) and (19) yields the estimate
$\biggl\|\sum_{k=1}^{8}y_{k}\lambda_{k}\biggr\|\;\leq\;C_{3}\,\max_{i,j}|\mu_{i}-\mu_{j}|\,\|X^{\dagger}X\|,$ (20)
which is precisely the bound stated in Eq. (15).
A.3 Remarks
The constant $C_{3}$ is not universal and depends on the chosen normalization of the generators. Its explicit numerical value is irrelevant for the formulation of Theorem 1. What matters is that $C_{3}$ is finite and depends only on the algebraic structure of $\mathfrak{su}(3)$, rather than on the specific map $\Phi$. This ensures that the condition (13) is purely algebraic and independent of the particular input operator $X$.